0.01 M solution of carbonic acid dissociates, and [HCO3-] = 0.000066M, then what is the Ka of carbonic acid?





1
Expert's answer
2020-10-14T06:02:35-0400

[HCO3-] = [H+] = 0,000066 M

[H2CO3] = C(H2CO3) - [H2CO3] = 0,01 - 0,000066 = 0,009934 M

Ka1, H2CO3 = "\frac{[HCO3^-][H^+]}{[H2CO3]} = \frac{[0,00066]^{2}}{[0,009934]} = 4,38 * 10^{-7}"